OpenStudy (anonymous):

for what integers will the sum of 10 times the integer and 4 times the reciprocal be 14?

OpenStudy (anonymous):

set up an equation: 10x+4(1/x)=14 then solve for x: (multiply each side by x to get rid of the x in the denominator) 10x^2+4=14x Then subtract 14x over: 10x^2-14x+4=0 Then FOIL and solve

OpenStudy (anonymous):

or the reverse of FOIL

OpenStudy (mertsj):

\[10x+4(\frac{1}{x})=14\] \[10x+\frac{4}{x}=14\] \[10x^2+4=14x\] \[10x^2-14x+4=0\] \[5x^2-7x+2=0\] \[(5x-2)(x-1)=0\] \[x=\frac{2}{5} , x=1\]

OpenStudy (anonymous):

\( 10x +4/x =14 \implies \left.\left\{x= \frac{2}{5}\right\},\{x= 1\}\right\} \) ony x=1 is the integer hence the answer.